OpenStudy (anonymous):

The hypotenuse of a right triangle is 37 mm. One leg of the right triangle is 12 mm. What is the length of the other leg? 30 mm 35 mm 40 mm 47 mm

OpenStudy (anonymous):

i got 25

OpenStudy (anonymous):

what is 37^2-12^2?

OpenStudy (anonymous):

nerd its 35

OpenStudy (anonymous):

don't give direct answers

OpenStudy (anonymous):

i still got 25

OpenStudy (anonymous):

ok so the sqr of 37 is 1369 take that and subtract 144 from it (sqr of 12) 1225 is the answer the sqr rt of 1225 is 35

OpenStudy (anonymous):

37^2=1369 12^2=144 1369-144=1225. sqrt tht

OpenStudy (anonymous):

1225

OpenStudy (anonymous):

exactly ty

OpenStudy (anonymous):

s0 1225 is 35?

OpenStudy (anonymous):

so*

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

1225 is \[35^2\]

OpenStudy (anonymous):

np

OpenStudy (arthur_ser):